OpenStudy (welshfella):

There are 5 cars red , blue , white , black and green and five keys each opening one of the cars. I pick a key and try to open one of the cars. What is the probability that it opens the blue car?

OpenStudy (calculusxy):

is this supposed to be a trick question?

OpenStudy (welshfella):

No

OpenStudy (calculusxy):

if not, then literally speaking i think it would be 1/5 = 20% is the probability to choose a key that will open the blue car.

OpenStudy (will.h):

My initial guess would be 1/5

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it has to open one of the cars right?

OpenStudy (misty1212):

if you had 6 keys to choose from, it would be different

OpenStudy (welshfella):

yes thats right. We are all pretty clever on OS! I's surprising how many people say its 1/5 * 1/5 = 1/25.

OpenStudy (calculusxy):

yay i am smart! thanks @welshfella for making me feel smart :)

OpenStudy (welshfella):

that's Ok lol

OpenStudy (will.h):

:) good day to you all.

OpenStudy (welshfella):

I think people just can't believe its so easy and over complicate it.

OpenStudy (calculusxy):

that's actually me most of the time ;)

OpenStudy (will.h):

I was surprised that you posted it anyway.

OpenStudy (calculusxy):

^^^^

OpenStudy (welshfella):

If the question was you have to pick the green key then the green car - that would be 1/25.

OpenStudy (calculusxy):

yes. that would be a combination problem right?

OpenStudy (will.h):

or since you will pick the green key at the 1st ty, then the green car will hav a doubled chance which means 2/5. or that's what i understand from your question

OpenStudy (welshfella):

No the 2 events are independent so you multiply the probabilities

Nnesha (nnesha):

Can i have the black car? ;;-;;

OpenStudy (welshfella):

yea OK

OpenStudy (welshfella):

I've been silly here. I got the question wrong . The way I wrote it the answer is 1/25 because you have to pick the blue key first and then pick the blue car. But if you say that there are 5 cars and 5 keys - one for each car . You pick any key then try and open one of the cars then the answer is 1/5. - Because it doesn't matter which key you pick in the first place. That was the original question . My only excuse is I'm tired!!